### v2

@@ 4.2 Multi-factor Case
Table 3: The estimation accuracy of factors when the number of factors $r=3$. The measure is defined as (28). $N$ and $T$ denote the dimension and the sample size, respectively. 500 replications are used in the experiments.
@@
$N$ 100 200 300 500 800 100 200 300 500 800
50 0.090 0.080 0.073 0.065 0.058 0.136 0.127 0.116 0.114 0.100
100 0.059 0.054 0.049 0.043 0.039 0.099 0.090 0.075 0.068 0.063
150 0.047 0.043 0.039 0.033 0.031 0.078 0.071 0.063 0.056 0.050
200 0.041 0.036 0.032 0.029 0.026 0.068 0.061 0.055 0.049 0.043
300 0.032 0.029 0.026 0.023 0.020 0.052 0.045 0.043 0.037 0.032
where $s=\left\lceil\sqrt{T}\right\rceil$ and $\widehat{S}_{j}$ is the the estimated nonzero locations in $\underline{\widehat{\mathbf{f}}}_{j}$ for $j=1,2$, and $3$. Table 4 reports the empirical accuracy of the sparsity locations in the factor processes when $r=3$. From Table 4, we see that the pattern is also similar to the case when $r=1$ in Section 4.1. The empirical results are in line with our asymptotic ones in the sense that the estimation accuracy will increase as the dimension or the sample size increases.
Table 4: Empirical recovery rate of the sparsity in the factor processes when the number of factors $r=3$. The error in each replication is calculated as (29). $N$ and $T$ denote the dimension and the sample size, respectively. 500 replications are used in the experiments.
@@
$N$ 100 200 300 500 800 100 200 300 500 800
50 0.949 0.950 0.953 0.957 0.960 0.931 0.931 0.932 0.934 0.940
100 0.966 0.965 0.965 0.969 0.971 0.947 0.948 0.952 0.955 0.958
150 0.971 0.971 0.973 0.977 0.976 0.956 0.957 0.960 0.964 0.965
200 0.973 0.974 0.977 0.977 0.980 0.959 0.959 0.964 0.966 0.970
300 0.980 0.979 0.981 0.983 0.983 0.968 0.971 0.971 0.973 0.975
Although the ratio-based method of (13) for determining the number of factors has been shown to be valid in many previous studies, we conduct an auxiliary experiment to verify its efficacy under our setting. Table 5 reports the empirical probabilities of $P(\widehat{r}=r)$ in 500 replications under the aforementioned setting. From Table 5, we see that the ratio-based method performs well, with most of the empirical probabilities being close to one. This is understandable since all the factors are strong ones. Similar results can be found in Gao and Tsay (2023).
Table 5: Empirical probabilities (EP) of determining the number of factors using the ratio-based method in (13) when the number of factors $r=3$. $N$ and $T$ denote the dimension and the sample size, respectively. 500 replications are used in the experiments.
@@
4.3 Determining the Sparsity with Cross-Validation
In this section, we study the estimation accuracy of the information criterion discussed in Section 2.6 and Remark 4 in estimating the sparsity of the factors. For simplicity, we only consider the case when $r=1$, but similar results can be obtained for $r>1$. The generation of the factors and the data is similar to those in Section 4.1, but we only keep the largest $s=\left\lceil\sqrt{T}\right\rceil$ elements of the factor process in absolute value for each sample size $T$. We consider the dimension $N=50,100,150,200$, and $300$, and the sample size $T=100,200,300,500$, and $800$ for each $N$ in this section. For each configuration of $(N,T)$, we set the number of partitions $J=1$ for simplicity and set $N_{1}=N_{2}=N/2$ in the partition. Since $s=\sqrt{T}$ is diverging with the sample size $T$, we will use the information criterion defined in (25) to estimate the sparsity. 500 replications are used throughout the experiment. Table 6 reports the Empirical probabilities (EP) of determining the sparsity parameter using the information criterion in (25) when the number of factors $r=1$. The empirical probabilities are calculated based on the 500 experiments. From Table 6, we see that the proposed information criterion works sufficiently well in estimating the sparsity parameters. When the idiosyncratic terms possess some dynamic dependence structure, the proposed method does not provide a probability close to one to estimate the sparsity, but the performance of the proposed method still improves as the sample size $T$ and/or the dimension $N$ increases, which is in agreement with our asymptotic theory in Theorem 4.
Table 6: Empirical probabilities (EP) of determining the sparsity parameter using the information criterion in (25) when the number of factors $r=1$. $N$ and $T$ denote the dimension and the sample size, respectively. 500 replications are used in the experiments.
@@ 4.3 Determining the Sparsity with Cross-Validation
200 1 1 1 1 1 1 1 1 1 1
300 1 1 1 1 1 1 1 1 1 1
5 An Empirical Application to Stock Returns
5.1 Data
Similar to Pelger (2020), we group these 332 stocks into 14 categories by industry sector, as shown in Table 7. The data encompasses a wide range of industry sectors within the stock market, including Oil, Finance, Electricity, Technology, Food, Manufacturing, Pharma&Chemicals, Primary Manufacturing, Machinery, Health, Transportation, Trade, Services, and Mining.
Table 7: Categorization of 332 Stocks in the S&P 500 Index by Industry Sector. The data spans from January 1, 2004, to December 31, 2016.
@@ 5.1 Data
5.2 Sparse Factor Estimation
We first determine the number of factors for the centered data ${\mathbf{X}}$ using the eigenvalue-ratio method in (13). Figure 2 plots the ratios of eigenvalues of ${\mathbf{X}}{\mathbf{X}}^{\prime}$ and shows that the largest gap between the eigenvalues occurs between $\widehat{\lambda}_{1}$ and $\widehat{\lambda}_{2}$, implying that the number of factors is $\widehat{r}=1$. Next, we apply Algorithm 1 and the cross-validation method in Section 2.6 with the number of random partitions $J=10$, where the seed number is set to be 1234 in R. For the sparsity of the factors over the time horizon, we conduct a grid search over the time interval $s\in[\lceil\sqrt{T}\rceil-10,\lceil\sqrt{T}\rceil+150]=[48,208]$, where $\lceil x\rceil$ denotes the smallest integer that exceeds $x$. The estimated sparsity parameter is $\widehat{s}=184$, indicating that there are 184 days during which the stocks were driven by significant systematic risks. We can see that the method achieves a substantial dimension-reduction compared to the full time length $T=3273$.
Figure 2: (a) The first $10$ eigenvalues of ${\mathbf{X}}{\mathbf{X}}^{\prime}$; (b) The plots of the ratios for the eigenvalues $\widehat{\lambda}_{i}$ of the matrix $\widehat{\mathbf{S}}$.
We then plot the estimated sparse factor returns in Figure 4. The plot reveals roughly three clusters of time periods associated with significant risks that systematically affect the stock market. The first cluster is centered around 2008, driven by the 2007–2008 financial crisis. The second and third clusters are closely related, both associated with the European sovereign debt crisis. The most prominent cluster is centered around 2008, driven by the 2007–2008 financial crisis. This period saw severe disruptions in the financial markets, with significant declines in stock returns across the board. For the two closely related clusters associated with the European sovereign debt crisis, the first cluster corresponds to the initial phase of the crisis in 2010, while the second cluster relates to the intensification of the crisis in 2011 and 2012.
Next, we estimate the loadings of the 332 stock returns given the sparse common factors. We present the estimated loadings, indicating the level of dependence of 14 different industry sectors on the common risk factor in Figure 3. The loadings quantify the extent to which each sector is influenced by these common factors. Here are the detailed observations based on the average dependence values: Oil Sector: Exhibits moderate dependence on common risk factors, with an average loading of 0.0122. Finance Sector: Shows significant dependence, with an average loading of 0.0163. Electricity Sector: Demonstrates relatively low dependence, with an average loading of 0.0061. Technology Sector: Reflects moderate dependence, with an average loading of 0.0094. Food Sector: Indicates low dependence on common factors, with an average loading of 0.0044. Manufacturing Sector: Displays moderate dependence, with an average loading of 0.0113. Pharma & Chemicals Sector: Shows relatively low dependence compared to other sectors, with an average loading of 0.0068. Primary Manufacturing Sector: Exhibits moderate dependence, with an average loading of 0.0130. Machinery Sector: Demonstrates moderate dependence, with an average loading of 0.0106. Health Sector: Indicates relatively low dependence, with an average loading of 0.0059. Transportation Sector: Shows moderate dependence, with an average loading of 0.0096. Trade Sector: Reflects moderate dependence, with an average loading of 0.0081. Services Sector: Exhibits moderate dependence, with an average loading of 0.0108. Mining Sector: Indicates relatively low dependence on common risk factors, with an average loading of 0.0077.
5.3 Bridging Time and Risk Factors
To further identify the systematic factors over time, referred to as time factors, Table LABEL:Table-ft in the Appendix provides a comprehensive list of dates with significant systematic risk factors, the reasons for these factors, and their associated time factors. Each entry explains the specific event and its impact on stock returns, summarized according to the descriptions of the time factors in Table 8. These reasons are extracted from the daily reports on CNN Money (www.money.cnn.com) after the market closes on each trading day.
From Table LABEL:Table-ft, we see that specific events impact the stock market on each date with identified systematic factors. We further analyze the words generated from the columns of "Reason" in Table LABEL:Table-ft through textual analysis. The word cloud shown in Figure 5 provides a visual representation of the most frequent terms and phrases mentioned as influencing factors. These words are associated with nine significant factors outlined in Table 8 that affect stock returns. We treat Credit risk, Oil price, China, and Europe as individual factors since they significantly influence stock returns during the period from January 1, 2004, to December 31, 2016. The nine factors are summarized in Table 8.
Table 8: The nine significant systematic risks over the the time horizon from January 1 2004 to December 31, 2016, identified by the proposed sparse factor modeling approach.
@@ 5.3 Bridging Time and Risk Factors
We also plot the frequency charts of the nine factors over the time horizon in Figure 6. The most frequently mentioned factor is market sentiment, highlighting its dominant role in driving stock price fluctuations. Economic indicators and government policies follow closely, indicating the significant impact of macroeconomic data and policy decisions on market behavior. Company-specific factors also play an important role, affecting individual stock performance and, by extension, broader market trends. Factors related to Europe, such as economic conditions and crises, have a moderate influence, reflecting the interconnectedness of global markets. Oil price fluctuations, China’s economic activities, and global events are also notable contributors, underscoring the importance of global economic dynamics. Credit risk, while mentioned less frequently, remains a critical factor during periods of financial instability.
Figure 5: Word cloud generated from the column of Reason in Table LABEL:Table-ft. The data consists of the daily stock returns for the balanced panel of S&P 500 stocks from January 1st 2004 to December 31st 2016.
Figure 6: Frequency of Time Factors Influencing Stock Market Movements from January 2004 to December 2016. The chart visualizes the occurrence of various time factors, including market sentiment, economic indicators, government policies, company-specific factors, Europe, oil price, China, global events, and credit risk, based on their frequency of mention in Table LABEL:Table-ft.
5.4 Implications
From Tables LABEL:Table-ft and 8, we can see that these factors are all important to the stability of the financial market. In general, they affect the financial market in the following ways:
@@ 5.4 Implications
From Tables LABEL:Table-ft and 8, we have the following findings and implications:
@@
6 Conclusion
Sparse factor modeling focuses on extracting meaningful insights by identifying a subset of relevant variables, thereby enhancing interpretability and reducing computational demands. This study proposes a novel perspective by assuming that factor processes are sparse over the timeline, rather than enforcing sparsity on loadings. This approach is particularly relevant for scenarios where systematic co-movements are significant only during certain periods, such as financial crises or policy changes.
The paper introduces approximate factor models and their estimation using conventional PCA, followed by the proposed sparse asymptotic PCA framework. It details the estimation procedures and theoretical properties. Simulation and empirical studies demonstrate the method’s effectiveness in providing interpretable results and linking co-movements to specific events. This study not only advances the methodology of sparse factor modeling but also bridges the gap between time-specific events and risk factors in economic and financial systems. The proposed approach offers a robust framework for future research, including the expansion of time horizons, sector-specific analyses, and the integration of advanced modeling techniques.
Supplementary Material for “Sparse Asymptotic PCA: Identifying Sparse Latent Factors Across Time Horizon"
A Examples of Sparse Factors in Time Horizons
  1. (i)
  2. (ii)
    Disaggregate Business Cycle Analysis. Gregory and Head (1999) found that cross-country variations have common components, referred to as global shocks. In addition, each country also experiences country-specific shocks. It is clear that the global and country-specific shocks are reflected by the common factors ${\mathbf{f}}_{t}$’s and the idiosyncratic terms ${\mathbf{e}}_{t}$’s in Model (2), respectively. Under the proposed framework, sparse factors over a certain time period correspond to situations where global shocks do not play a dominant role in driving the cyclical variations in a country’s economy. Conversely, nonzero factors will inspire researchers to discover the possible reasons for significant global shocks during the corresponding time period.
  3. (iii)
  4. (iv)
    Monitoring and Forecasting. Stock and Watson (1998) and Forni et al. (2000b) have demonstrated that factor models provide an effective framework for monitoring economic activities, as large-dimensional economic panel data can often be summarized by a small number of factors. In this context, business cycles can be modeled through the co-movements of economic variables, which correspond to the common factors ${\mathbf{f}}_{t}$ in Model (2). When these factors exhibit sparsity, our approach allows us to track economic activities by identifying systematic risks over time. Furthermore, periods of heightened volatility in the economic panel may produce nonzero factors that are associated with specific events that influence the overall system.
B Mathematical Proofs
We will use $C$ or $c$ to denote a generic constant the value of which may change at different places. We first introduce several notations used in the derivations. Let ${\mathbf{V}}={\mathbf{F}}/\sqrt{T}$, and ${\mathbf{v}}_{i}=\underline{{\mathbf{f}}}_{i}/\sqrt{T}$ be the $i$-th columns of ${\mathbf{V}}$. Define ${\mathbf{S}}_{c}={{\mathbf{F}}\boldsymbol{\Lambda}^{\prime}\boldsymbol{\Lambda}{\mathbf{F}}^{\prime}}/{(NT)}$ and ${\mathbf{S}}_{e}={\mathbf{e}}{\mathbf{e}}^{\prime}/(NT)$. It follows from Model (3) and Assumptions 24 that
@@ B Mathematical Proofs
where ${\mathbf{S}}_{R}={\mathbf{S}}_{e}+\boldsymbol{\Sigma}_{fe}+\boldsymbol{\Sigma}_{fe}^{\prime}$ and $\boldsymbol{\Sigma}_{fe}={\mathbf{F}}\boldsymbol{\Lambda}^{\prime}{\mathbf{e}}^{\prime}/(NT)$ We need a few lemmas first.
@@
For the $i$-th iteration in Algorithm 2 of the main article, the matrix ${\mathbf{B}}_{i}$ in step 7 of Algorithm 2 is symmetric for each $i=1,...,r+1$.
Proof. We will prove the result by an induction method. For each iteration $i=1,...,r+1$, we will first show that
@@
where we used the assumption in (S.2) for $i=j$ and the fact that $\widehat{\mathbf{v}}_{j}^{\prime}{\mathbf{B}}_{j}\widehat{\mathbf{v}}_{j}=1$. Therefore, this proves (S.2) for all $i$. It follows from (S.2) that ${\mathbf{B}}_{i}$ is a symmetric matrix for all $i=1,...,r+1$. This completes the proof. $\Box$
@@
where there are $T-2$ zeros.
@@
where there are $T-2$ zeros. This completes the proof. $\Box$
@@
Under Assumptions 17, the estimator $\widehat{\mathbf{v}}_{1}$ of ${\mathbf{v}}_{1}$ satisfies
@@
By (B) and (S.6), we replace ${\mathbf{v}}$ by the estimator $\widehat{\mathbf{v}}_{1}$ of ${\mathbf{v}}_{1}$ and obtain
@@
$\mbox{tr}({\mathbf{S}}(\widehat{\mathbf{v}}_{1}\widehat{\mathbf{v}}_{1}^{\prime}-{\mathbf{v}}_{1}{\mathbf{v}}_{1}^{\prime}))=\widehat{\mathbf{v}}_{1}^{\prime}{\mathbf{S}}\widehat{\mathbf{v}}_{1}-{\mathbf{v}}_{1}^{\prime}{\mathbf{S}}{\mathbf{v}}_{1}>0$ (S.8)
@@
Let Assumptions 17 hold. For any ${\mathbf{v}}\in S_{1}$, we have
@@
Let Assumptions 17 hold. For any vector ${\mathbf{v}}\in S_{1}$, we have
@@
where $Z_{j}$ is defined as that in the proof of Lemma A.6. Note that $W_{j}$ and $Z_{j}$ are independent of each other and they all have finite variances. In addition, Assumption 5 implies that $Z_{j}$ are independent across $j=1,...,N$. By Assumption 7, $W_{j}Z_{j}$ satisfies condition (2.7) in Merlevède et al. (2011) with $\gamma_{2}=1$. Since $Z_{j}$ are independent over $j=1,...,N$, then condition (2.6) in Merlevède et al. (2011) is satisfied for any $\gamma_{1}>0$. By Theorem 1 of Merlevède et al. (2011), we have
@@
Therefore, we can obtain from (B) that
@@
Proof of Theorem 2. According to Algorithms 1–3, we can repeat the proofs of Theorem 1 for the eigenvectors obtained in the subsequent iterations. So long as the previous eigenvector is consistent, we can similarly obtain the consistencies of the next few estimated eigenvectors. On the other hand, because the number of factors $r$ is finite, by (3), we have
$\rho(\widehat{\mathbf{V}}_{1},\widehat{\mathbf{V}}_{1})^{2}=2\mbox{tr}\left({\mathbf{I}}_{r}-\widehat{\mathbf{V}}^{\prime}{\mathbf{V}}{\mathbf{V}}^{\prime}\widehat{\mathbf{V}}\right).$
Note that the columns in the true factor matrix ${\mathbf{F}}$ are orthogonal, then the cross terms in $\widehat{\mathbf{V}}{\mathbf{V}}$ are of the same order as that in Theorem 1. We can apply the previous results to each component of the one on the right-hand side, and the result in Theorem 2 holds since $r$ is finite. See also the descriptions for different distances between two matrices in Equations (18)–(3) of the main text. This completes the proof. $\Box$
@@
Proof of Theorem 3. Let $\widehat{\mathbf{F}}$ be the estimated sparse factors by the proposed algorithms. We apply the least-squares method to (3) and obtain
@@
where ${\mathbf{H}}_{s}=(\widehat{\mathbf{F}}^{\prime}\widehat{\mathbf{F}})^{-1}\widehat{\mathbf{F}}^{\prime}{\mathbf{F}}$. By Assumption 7 and the results in Theorem 1, it is not hard to show that
$\max_{1\leq i\leq N}|\frac{\widehat{\mathbf{F}}^{\prime}{\mathbf{e}}_{i}}{T}|$ $\leq\max_{1\leq i\leq N}|\frac{1}{T}(\widehat{\mathbf{F}}-{\mathbf{F}})^{\prime}{\mathbf{e}}_{i}|+\max_{1\leq i\leq N}|\frac{1}{T}{\mathbf{F}}^{\prime}{\mathbf{e}}_{i}|$
$\leq=O_{p}(\sqrt{\frac{s^{*}{{}^{2}}\log(N)}{NT^{2}}}\log(T)+\sqrt{\frac{\log(N)}{T}}).$
Since $s^{*}=T^{\alpha}$ where $\alpha=\max(\alpha_{1},...,\alpha_{r})$ and $\alpha_{i}$ is defined in Assumption 3, if $\sqrt{\frac{1}{NT^{1-2\alpha}}}\log(T)=o(1)$, then it reduces to
$\max_{1\leq i\leq N}|\frac{\widehat{\mathbf{F}}^{\prime}{\mathbf{e}}_{i}}{T}|=O_{p}(\sqrt{\frac{\log(N)}{T}}).$
Then Theorem 3(i) holds.
@@
Note that $\frac{\widehat{\mathbf{F}}^{\prime}\widehat{\mathbf{F}}}{T}\rightarrow_{p}{\mathbf{I}}_{r}$ due to the average consistency of each estimated factor process and Assumption 2, by Assumptions 56 and the results in Theorem 2, we can show that
@@
Let Assumptions 17 hold. For any $s$ with $1\leq s\leq s_{0}$, we have

### v3

@@ B Some Tables and Figures for Numerical Studies
Table S.I: The estimation accuracy of factors when the number of factors $r=3$. The measure is defined as (4.3) of the main text. NN and $T$ denote the dimension and the sample size, respectively. 500 replications are used in the experiments.
@@
NN 100 200 300 500 800 100 200 300 500 800
50 0.090 0.096 0.097 0.099 0.010 0.136 0.159 0.158 0.158 0.155
100 0.059 0.064 0.065 0.066 0.067 0.099 0.105 0.105 0.105 0.112
150 0.047 0.050 0.052 0.053 0.053 0.078 0.087 0.086 0.085 0.086
200 0.041 0.043 0.044 0.045 0.045 0.068 0.070 0.073 0.073 0.072
300 0.032 0.034 0.034 0.035 0.036 0.052 0.053 0.057 0.058 0.057
Table S.II: Empirical recovery rate of the sparsity in the factor processes when the number of factors $r=3$. The error in each replication is calculated as (4.4) of the main text. NN and $T$ denote the dimension and the sample size, respectively. 500 replications are used in the experiments.
@@
NN 100 200 300 500 800 100 200 300 500 800
50 0.949 0.946 0.945 0.942 0.941 0.931 0.931 0.918 0.918 0.917
100 0.966 0.960 0.960 0.958 0.957 0.947 0.941 0.941 0.940 0.937
150 0.971 0.967 0.966 0.964 0.964 0.956 0.949 0.949 0.949 0.948
200 0.973 0.971 0.969 0.970 0.969 0.959 0.958 0.953 0.955 0.955
300 0.980 0.974 0.977 0.975 0.974 0.968 0.964 0.965 0.963 0.962
Table S.III: Empirical probabilities (EP) of determining the number of factors using the ratio-based method in (2.12) of the main text when the number of factors $r=3$. NN and $T$ denote the dimension and the sample size, respectively. 500 replications are used in the experiments.
@@
Table S.IV: Empirical probabilities (EP) of determining the sparsity parameter using the information criterion in (3.9) when the number of factors $r=1$. NN and $T$ denote the dimension and the sample size, respectively. 500 replications are used in the experiments.
@@
Table S.V: Categorization of 332 Stocks in the S&P 500 Index by Industry Sector. The data spans from January 1, 2004, to December 31, 2016.
@@
Figure S.3: (a) The first $10$ eigenvalues of ${\mathbf{X}}{\mathbf{X}}^{\prime}$; (b) The plots of the ratios for the eigenvalues $\widehat{\lambda}_{i}$ of the matrix $\widehat{\mathbf{S}}$.
Table S.VI: The nine significant systematic risks over the time horizon from January 1 2004 to December 31, 2016, identified by the proposed sparse factor modeling approach.
@@
B.1 Further Implications of the Empirical Analysis
From Tables LABEL:Table-ft and S.VI, we can see that these factors are all important to the stability of the financial market. In general, they affect the financial market in the following ways:
@@ B.1 Further Implications of the Empirical Analysis
From Tables LABEL:Table-ft and S.VI, we have the following findings and implications:
@@
C Mathematical Proofs
We will use $C$ or $c$ to denote a generic constant the value of which may change at different places. We first introduce several notations used in the derivations. Let ${\mathbf{V}}={\mathbf{F}}/\sqrt{T}$, and ${\mathbf{v}}_{i}=\underline{{\mathbf{f}}}_{i}/\sqrt{T}$ be the $i$-th columns of $\mathbf{V}$}. Define ${\mathbf{S}}_{c}={{\mathbf{F}}\boldsymbol{\Lambda}^{\prime}\boldsymbol{\Lambda}{\mathbf{F}}^{\prime}}/{(NT)}$ and ${\mathbf{S}}_{e}={\mathbf{e}}{\mathbf{e}}^{\prime}/(NT)$. It follows from Model (3) and Assumptions 2–4 of the main article that
@@ C Mathematical Proofs
where ${\mathbf{S}}_{R}={\mathbf{S}}_{e}+\boldsymbol{\Sigma}_{fe}+\boldsymbol{\Sigma}_{fe}^{\prime}$ and $\boldsymbol{\Sigma}_{fe}={\mathbf{F}}\boldsymbol{\Lambda}^{\prime}{\mathbf{e}}^{\prime}/(NT)$ . We need a few lemmas first.
@@
For the $i$-th iteration in Algorithm 2 of the main article, the matrix ${\mathbf{B}}_{i}$ in step 7 of Algorithm 2 is symmetric for each i=1,,r+1i=1,...,r+1.
Proof. We will prove the result by an induction method. For each iteration i=1,,r+1i=1,...,r+1, we will first show that
@@
where we used the assumption in (S.2) for $i=j$ and the fact that $\widehat{\mathbf{v}}_{j}^{\prime}{\mathbf{B}}_{j}\widehat{\mathbf{v}}_{j}=1$. Therefore, this proves (S.2) for all $i$. It follows from (S.2) that ${\mathbf{B}}_{i}$ is a symmetric matrix for all i=1,,r+1i=1,...,r+1. This completes the proof. $\Box$
@@
where there are T2T-2 zeros.
@@
where there are T2T-2 zeros. This completes the proof. $\Box$
@@
Under Assumptions 1–7, the estimator $\widehat{\mathbf{v}}_{1}$ of ${\mathbf{v}}_{1}$ satisfies
@@
By (C) and (S.6), we replace $\mathbf{v}$} by the estimator $\widehat{\mathbf{v}}_{1}$ of ${\mathbf{v}}_{1}$ and obtain
@@
$\mbox{tr}({\mathbf{S}}(\widehat{\mathbf{v}}_{1}\widehat{\mathbf{v}}_{1}^{\prime}-{\mathbf{v}}_{1}{\mathbf{v}}_{1}^{\prime}))=\widehat{\mathbf{v}}_{1}^{\prime}{\mathbf{S}}\widehat{\mathbf{v}}_{1}-{\mathbf{v}}_{1}^{\prime}{\mathbf{S}}{\mathbf{v}}_{1}>0.$ (S.8)
@@
Let Assumptions 1–7 hold. For any ${\mathbf{v}}\in S_{1}$, we have
@@
Let Assumptions 1–7 hold. For any vector ${\mathbf{v}}\in S_{1}$, we have
@@
where $Z_{j}$ is defined as that in the proof of Lemma A.6. Note that $W_{j}$ and $Z_{j}$ are independent of each other and they all have finite variances. In addition, Assumption 5 implies that $Z_{j}$ are independent across j=1,,Nj=1,...,N. By Assumption 7, $W_{j}Z_{j}$ satisfies condition (2.7) in Merlevède et al. (2011) with $\gamma_{2}=1$. Since $Z_{j}$ are independent over j=1,,Nj=1,...,N, then condition (2.6) in Merlevède et al. (2011) is satisfied for any $\gamma_{1}>0$. By Theorem 1 of Merlevède et al. (2011), we have
@@
Therefore, we can obtain from (C) that
@@
Proof of Theorem 2. According to Algorithms 1–3, we can repeat the proofs of Theorem 1 for the eigenvectors obtained in the subsequent iterations. So long as the previous eigenvector is consistent, we can similarly obtain the consistencies of the next few estimated eigenvectors. On the other hand, because the number of factors $r$ is finite, by (3.5), we have
[…]
Note that the columns in the true factor matrix $\mathbf{F}$} are orthogonal, then the cross terms in $\widehat{\mathbf{V}}{\mathbf{V}}$ are of the same order as that in Theorem 1 of the main text. We can apply the previous results to each component of the one on the right-hand side, and the result in Theorem 2 holds since $r$ is finite. See also the descriptions for different distances between two matrices in Equations (3.2)–(3.5) of the main text. This completes the proof. $\Box$
@@
Proof of Theorem 3. Let $\widehat{\mathbf{F}}$ be the estimated sparse factors by the proposed algorithms. We apply the least-squares method to (2.3) of the main text and obtain
@@
where ${\mathbf{H}}_{s}=(\widehat{\mathbf{F}}^{\prime}\widehat{\mathbf{F}})^{-1}\widehat{\mathbf{F}}^{\prime}{\mathbf{F}}$. Note that
$\max_{1\leq i\leq N}\left|\frac{\widehat{{\mathbf{F}}}^{\prime}{\mathbf{e}}_{i}}{T}\right|$ $\leq\max_{1\leq i\leq N}\left|\frac{1}{T}(\widehat{{\mathbf{F}}}-{\mathbf{F}})^{\prime}{\mathbf{e}}_{i}\right|+\max_{1\leq i\leq N}\left|\frac{1}{T}{\mathbf{F}}^{\prime}{\mathbf{e}}_{i}\right|.$
We now consider the first term. Let $\widetilde{{\mathbf{F}}}$ denote the PCA factors obtained without thresholding. Since (S.18) is a least-squares method, we construct the sparse factor estimates $\widehat{{\mathbf{F}}}$ by retaining only the largest $\mathbf{s}$} elements (in absolute value) of $\widetilde{{\mathbf{F}}}$ columnwise, using our algorithm without normalization.
Without loss of generality, suppose
${\mathbf{F}}=\begin{pmatrix}{\mathbf{F}}_{1}\\ {\mathbf{F}}_{2}\end{pmatrix},\quad\widetilde{{\mathbf{F}}}=\begin{pmatrix}\widetilde{{\mathbf{F}}}_{1}\\ \widetilde{{\mathbf{F}}}_{2}\end{pmatrix},\quad\widehat{{\mathbf{F}}}=\begin{pmatrix}\widehat{{\mathbf{F}}}_{1}\\ \widehat{{\mathbf{F}}}_{2}\end{pmatrix},\quad{\mathbf{e}}_{i}=\begin{pmatrix}{\mathbf{e}}_{i,1}\\ {\mathbf{e}}_{i,2}\end{pmatrix}$,
where ${\mathbf{F}}_{1}$ contains the non-sparse components and ${\mathbf{F}}_{2}=\mathbf{0}$. Under the thresholding procedure, $\widehat{{\mathbf{F}}}_{1}=\widetilde{{\mathbf{F}}}_{1}$ and $\widehat{{\mathbf{F}}}_{2}=\mathbf{0}$. Then,
$\max_{1\leq i\leq N}\left|\frac{1}{T}(\widehat{{\mathbf{F}}}-{\mathbf{F}})^{\prime}{\mathbf{e}}_{i}\right|$ $\leq\max_{1\leq i\leq N}\left|\frac{1}{T}(\widehat{{\mathbf{F}}}-\widetilde{{\mathbf{F}}})^{\prime}{\mathbf{e}}_{i}\right|+\max_{1\leq i\leq N}\left|\frac{1}{T}(\widetilde{{\mathbf{F}}}-{\mathbf{F}})^{\prime}{\mathbf{e}}_{i}\right|$
$\leq\max_{1\leq i\leq N}\left|\frac{1}{T}\widetilde{{\mathbf{F}}}_{2}^{\prime}{\mathbf{e}}_{i,2}\right|+\max_{1\leq i\leq N}\left|\frac{1}{T}(\widetilde{{\mathbf{F}}}-{\mathbf{F}})^{\prime}{\mathbf{e}}_{i}\right|.$
By the argument in Lemma B.1 of Bai (2003), the second term is at most of order $\delta_{NT}^{-2}=O_{p}(1/N+1/T^{2})$ for each $i$, where we have used the improved rate $\delta_{NT}^{-1}=O_{p}(\sqrt{1/N}+1/T)$ as established in Proposition 1 of Bai and Ng (2023). The first term can be handled similarly by partitioning the true factors and the PCA-estimated factors, following the proof of Lemma B.1 in the same reference. Therefore, the first term is also at most of order $O_{p}(1/N+1/T^{2})$ for each $i$. Consequently,
$\max_{1\leq i\leq N}\left|\frac{\widehat{{\mathbf{F}}}^{\prime}{\mathbf{e}}_{i}}{T}\right|$ $=O_{p}\left(\left(\frac{1}{N}+\frac{1}{T^{2}}\right)\sqrt{\log N}\right)+O_{p}\left(\sqrt{\frac{\log N}{T}}\right)=O_{p}(\frac{\sqrt{\log N}}{N}+\sqrt{\frac{\log N}{T}}).$
If $\sqrt{T}/N=o(1)$, then the bound simplifies to
$\max_{1\leq i\leq N}\left|\frac{\widehat{{\mathbf{F}}}^{\prime}{\mathbf{e}}_{i}}{T}\right|=O_{p}\left(\sqrt{\frac{\log N}{T}}\right).$
This completes the verification of Theorem 3(i).
@@
Note that $\frac{\widehat{\mathbf{F}}^{\prime}\widehat{\mathbf{F}}}{T}\rightarrow_{p}{\mathbf{I}}_{r}$ due to the average consistency of each estimated factor process and Assumption 2, by Assumptions 5–6 and the results in Theorem 2, we can show that
@@
Proof of Corollary 1. We first show the consistency of the information criterion in (2.11). Since the upper bound of the estimated factors is $C_{NT}=\sqrt{\log(T)/N}+1/T$, it suffices to verify the conditions in Theorem 2 of Bai and Ng (2002):
$\frac{N+T}{NT}\log(T)\log\left(\frac{NT}{N+T}\right)\rightarrow 0,\quad\text{and}\quad C_{NT}^{-2}\frac{N+T}{NT}\log(T)\log\left(\frac{NT}{N+T}\right)\rightarrow\infty.$ (S.21)
These hold straightforwardly under the given $C_{NT}$, ensuring that the conditions in Bai and Ng (2002) are met. Hence, the information criterion in (2.11) consistently estimates $r$.
Next, we establish the consistency of the eigenvalue ratio method in (2.12). According to Equations (2)–(3) and the subsequent discussion in Ahn and Horenstein (2013), the conditions stated in Corollary 1, along with Assumptions 1–7, satisfy all the requirements of Theorem 1 in Ahn and Horenstein (2013). Therefore, the ratio-based estimator in (2.12) also consistently estimates $r$. This completes the proof of the corollary. $\Box$
To prove Theorem 4, we need the following lemmas.
@@
Let Assumptions 1–7 hold. For any $s$ with $1\leq s\leq s_{0}$, we have